OpenStudy (susanna):

Write an equation of the tangent function with period pi/4, phase shift pi, and vertical shift 1.

OpenStudy (susanna):

@SolomonZelman

OpenStudy (michele_laino):

hint: a tangent function with a period of pi/4, if can be written as follows: \(y=\tan(4x)\)

OpenStudy (susanna):

I chose this answer...am I right? y=tan(4x-4pi)+1

OpenStudy (michele_laino):

I think it is correct!

OpenStudy (michele_laino):

since we can rewrite such function, as below: \(y=\tan(4(x-\pi))+1\) and in that form is evident the phase shift of \(\pi\) radians